Find the Maclaurin series of Tore ed dt. Σ n=0
Answer #1

Given ** x we know that e* = 1+ 1! + + + + 2! 3! 4! 2 = 2-0 ! replace x by - (-20) II IM IME -1)*26* = n! X (-1)* <*+2 0 ***